Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-181_2001-187

Mission Operations Days: 2001/181 to 2001/187
June 30, 2001 00:00:00 GMT to July 06, 2001 00:00:00 GMT


MODIS powered up on PS1/CPA, on day 183, 20:59:00, and is now collecting science data (see below).

TMON 69 built and loaded, but not currently enabled. Monitors EPS current for rise above 1.8 amps. If so, MODIS shuts down.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On; #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
Nominal
Inhibit Ids SetNoneNominal
TMON 66 (SRCA A)DisabledNominal
TMON 67 (SRCA B)DisabledNominal
TMON 68 (FIFO)EnabledNominal
TMON 69 (EPS Current)DisabledNominal

This Week's Completed MODIS Activities:

Saturday, June 30, 2001

None

Sunday, July 1, 2001

None

Monday, July 2, 2001

2001/183 19:06:50 – 19:09:04 Real Time Turned on MODIS using power supply 1 (PS1), control processor B (CPB).
2001/183 19:14:32 – 19:14:51 Real Time Reset Control Processor B to Upload Mode
2001/183 19:16:31 – 19:17:17 Real Time Dumped CPB Startup code memory
2001/183 19:39:41 – 19:40:01 Real Time Restarted CPB with ASTATE = 0
2001/183 20:06:04 – 20:10:43 Real Time Turned off MODIS
2001/183 20:57:04 – 20:58:52 Real Time Turned on MODIS using PS1, CPA.
2001/183 21:14:24 – 21:15:05 Real Time Ran macro 0 to clear spurious loads
2001/183 22:48:16 Real Time MODIS transition to A-side SCIENCE mode
2001/183 22:52:29 Real Time Reset Formatter A to Upload Mode
2001/183 22:53:20 – 22:54:21 Real Time Loaded two Formatter Anomaly recovery memory patches
2001/183 22:55:02 Real Time Restarted formatter A with ASTATE = 0

Tuesday, July 3, 2001

2001/184 00:19:20 – 00:21:11 Real Time Dumped formatter error counts and addresses
2001/184 00:21:35 – 00:24:39 Real Time Enabled telemetry alarms for formatter anomaly mnemonics and disabled blackbody temperature alarms
2001/184 00:25:20 – 00:35:31 Real Time Dump formatter anomaly patches to verify integrity
2001/184 18:06:50 – 18:08:10 Real Time Turned on A-side Black Body (BB) to 315K
2001/184 18:08:47 – 18:09:31 Real Time Set Blackbody temperature to 290K
2001/184 18:10:21 – 18:10:27 Real Time Enabled Blackbody temperature alarms

Wednesday, July 4, 2001

None

Thursday, July 5, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

In the month of July, MODIS will be performing SRCA spatial, radiometric, and spectral calibrations, a black body calibration, and weekly SD/SDSM calibrations.

Friday, July 6, 2001

187/20:55:17 ATC OA-16: SD/SDSM Screened Calibration
187/22:34:10 ATC OA-15: SD/SDSM Open Calibration

Saturday, July 7, 2001

None

Sunday, July 8, 2001

None

Monday, July 9, 2001

None

Tuesday, July 10, 2001

191/16:45? Real-time FOT Set IRU to high rate for roll maneuver; less pointing accuracy
191/18:09:37 ATC Move Formatter to Night ahead by 3:23 to ensure 50/50 day/night
191/18:18:53 ATC FOT begins spacecraft roll slew to -11.7504 degrees
191/18:24:40 ATC Set Formatter to Day Rate
191/18:24:42 ATC Set SCIABNORM Flag to ABNORM
191/18:24:44 ATC Set PC band DC Restore to OFF
191/18:24:46 ATC Set PV band DC Restore to OFF
191/18:24:48 ATC Set Sector Rotation to -3072 (Earth Aperture to view Space view and OBCs)
191/18:31:18 ATC Set Sector Rotation to Zero (Normal)
191/18:31:20 ATC Set PV band DC Restore to ON
191/18:31:22 ATC Set PC band DC Restore to ON
191/18:31:24 ATC Set SCIABNORM Flag to NORM
191/18:31:26 ATC Set Formatter to Night Rate
191/18:36:33 ATC Spacecraft has returned to nominal attitude
191/18:45? Real-time FOT set IRU to low rate after roll maneuver; nominal pointing
191/19:03:51 ATC Move Formatter to Day later by 1.23 to ensure 50/50 day/night
Note: should have been 19:05:51, collecting 2 extra minutes of day rate data

Wednesday, July 11, 2001

192/10:33:05 - 11:30:36 ATC OA-23: SRCA Full Spatial
192/21:14:00 ATC OA-15: SD/SDSM Open
192/22:05:18 - 22:38:51 ATC OA-19: SRCA Full Radiometric
192/22:52:53 ATC OA-16: SD/SDSM Screened

Thursday, July 12, 2001

193/12:53:15 - 13:44:54 ATC OA-22: SRCA Full Spectral Part 1A
193/14:47:34 - 15:22:19 ATC OA-22: SRCA Full Spectral Part 2A
193/16:11:10 - 17:02:26 ATC OA-22: SRCA Full Spectral Part 3A
193/18:05:21 - 18:43:19 ATC OA-22: SRCA Full Spectral Part 4A
193/21:00:00 ATC Set Blackbody temperature to 270K

Friday, July 13, 2001

194/15:00:00 ATC Set Blackbody temperature to 280K
194/19:00:00 ATC Set Blackbody temperature to 285K
194/22:00:00 ATC Set Blackbody temperature to 290K

Saturday, July 14, 2001

195/02:00:00 ATC Set Blackbody temperature to 295K
195/02:00:30 ATC Set Blackbody duty cycle to FULL
195/04:00:00 ATC Set Blackbody temperature to 300K
195/06:00:00 ATC Set Blackbody temperature to 315K
195/09:00:00 ATC Set Blackbody temperature to 270K
195/09:00:30 ATC Set Blackbody duty cycle to THIRD

Sunday, July 15, 2001

196/05:00:00 ATC Set Blackbody temperature to 290K

Maneuvers:

The next MODIS Lunar Roll Calibration is scheduled for DOY 191 from 18:24:48 to 18:31:18. The roll was delayed by two orbits to avoid conflict with the CLAMS field campaign.

MODIS Anomalies:


Quick Summary:

On July 2, 2001 (DOY 183), at 20:59:00, MODIS was powered up on power supply 1 (PS1), control processor A (CPA). It was then commanded to A-side SCIENCE mode. The events leading up to this are as follows:


On June 15th, 2001, at 03:56:08 GMT, MODIS experienced an anomaly.

The Flight Operations Team paged the on-call MODIS engineer reporting a trip of TMON #68 (autonomous Telemetry Monitoring software on the spacecraft). TMON #68 is the MODIS double FIFO (First In, First Out data buffer) Write State TMON. In addition, many red limit violations were reported. Inspection of telemetry showed peculiar states including a CP (Control Processor) mode of INVALID and the MODIS model as ENGINEERING. With all telemetry suspect, the MODIS IOT commanded the instrument to low-power mode (using procedure MOD_ALL_OFF) at 05:16:43.

Evidence was very strong that a failure had occurred in PS2 and that the rest of the instrument was okay. There was still the question about whether the failure was a soft failure or a hard failure. If the failure was a hard failure (such as an electrical short), it was possible that powering up PS2 could possibly stress the loads (those components still internally enabled on MODIS from the time of the PS2 shut off) with an over voltage. However, if the failure was a soft failure (such as a Single Event Upset (SEU) caused by a high-energy particle impact), it was possible that powering up PS2 would return it to full operational capability.

On June 28th (day of year 179), at 18:59:30, MODIS was powered on using power supply 1 (PS1) and control processor B (CPB). The 1553 bus stopped sending telemetry, and at the time, there were questions about the current values. At 19:09:00, the MOD_ALL_OFF procedure was executed and MODIS was again turned OFF.

On July 2nd (day of year 183), at 19:06:50, MODIS was once again powered ON using PS1 and CPB. After commanding to upload mode and doing a formatter restart, everything appeared to be okay for the remainder of the contact. At the beginning of the next contact,, MODIS was no longer sending valid telemetry (master cycle was no longer incrementing). At this time, MODIS was powered OFF. At 20:57:04, MODIS was powered up using PS1 and CPA. Since everything looked nominal at this time, MACRO 0 was executed in real time by the MODIS IOT. At 22:48:16, MODIS was transitioned to A-side SCIENCE mode, and software patches were loaded to account for the formatter A anomaly. At the time of the transition to side A science, the formatter reset itself 65536 times in the first 4 hours 40 minutes (half of which came in the first 5 minutes). However, since approximately 2 PM on July 3rd, no new formatter resets have been received.

As of this time, MODIS is operating nominally on PS1, CPA. L1A science data is also nominal.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Voltage, current and temperature trending over the previous month, in addition to selected past time periods is currently underway to support the ongoing analysis of the PS2 shutdown anomaly.

Non-MODIS Significant Events:

SSR Anomaly:

Late in the evening of June 16, 2001, the Solid State Recorder (SSR) locked up. The SSR was no longer recording data and was not responding to commands uplinked by the Terra Flight Operations Team (FOT) to dump its data. This made it impossible for the ground to acquire any recorded housekeeping telemetry or science data from the Terra spacecraft.

After further testing and analysis, the SSR was restored to full operations (successfully dumping science data) at 11:40:34 GMT on June 20, 2001.

An interesting byproduct of this event was the functional restoration of supersets 30 and 31 (which, after reallocation of the supersets, they now reside in the MISR buffer). Additionally, ASTER's superset allocation was reduced by one superset to solve a complication with replays of their science data. This free superset has been allocated to the MODIS buffer. With the restoration of the malfunctioning supersets (which are no long allocated to the MODIS buffer), and the additional superset, MODIS will have approximately 3% greater science data storage space than originally allocated.

Limited Life Item Status:

SRCA 10W Lamp #1: 186.1 of 500 hours
SRCA 10W Lamp #2: 143.0 of 500 hours
SRCA 10W Lamp #3: 152.2 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.6 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1633 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements